gleaskad:

Find the quotient. (5ab - 10b 2 + 15bc) ÷ 5b

Eiwoh2:

Well, how would you simplify this?

Nnesha:

\[\rm \rm \frac{ 5ab-10b^2+15bc }{ 5b }\] rewrite as\[\rm \frac{ 5ab }{ 5b }-\frac{ 10b^2 }{ 5b }+\frac{\rm 15bc }{ 5b}\]